the height of a cylinder measures 4 cm more than its radius. find the height of the cylinder if its total area is 140pi?
Answer by stanbon(75887)

the height of a cylinder measures 4 cm more than its radius. find the height of the cylinder if its total area is 140pi?
---
Total surface area = 2(pi)r^2 + 2(pi)r*h
Solve:
140 = 2(pi)r^2 + 2(pi)*r(r+4)
70 = (pi)r^2 + (pi)(r^2+4r)
70 = 2(pi)r^2 + 4(pi)r
35 = (pi)r^2 + 2(pi)r
Rearrange:
(pi)r^2 + 2(pi)r - 35 = 0
r = 2.484 cm
h = 6.484 cm
====================
Cheers,
Stan H.
===================================